for problems 7-9 solve the system of equations using addition to eliminate one variable. if the system has no solution indicate that it is inconsistent. if it has infinite solutions, indicate that it is dependent. if the system check your answer.
3x-2y=-14
3x+5y=-14
check if possible

Answer by dabanfield(803)

3x-2y=-14
3x+5y=-14
Subtract the terms of the two equations:
(3-3)*x + (-2-5)*y = -14 - (-14)
0*x -7*y = 0
-7y = 0
y = 0
Substitute y=0 in the first equation:
3*x - 2*0 = -14
3*x = -14
x = -14/3